OpenStudy (13ziar):

PLEASE turn the following into symbol equations: potassium chloride + bromine potassium bromine + iodine potassium chloride + iodine

OpenStudy (13ziar):

@zaynab123 @Somy

OpenStudy (somy):

first off, lets start by simply writing their symbols. Do it and show me

OpenStudy (13ziar):

potassium = K chlorine/chloride = C/C2 bromine/bromide = Br/Br2

OpenStudy (somy):

good so now write it as they ask you to

OpenStudy (13ziar):

potassium chloride + bromine = KCl2+Br potassium bromine + iodine = KBr+I2 potassium chloride + iodine = KCl+I2

OpenStudy (13ziar):

is that correct

OpenStudy (somy):

potassium chloride + iodine = KCl+I2 - correct potassium bromine + iodine = KBr+I2 - correct potassium chloride + bromine = KCl2+Br - incorrect

OpenStudy (13ziar):

oh so would it be KCL+Br2

OpenStudy (somy):

yep

OpenStudy (13ziar):

thanks